OpenStudy (anonymous):

What is the value of f[g(6)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5? What is the value of f[g(6)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5? @Mathematics

OpenStudy (ashish56):

15

OpenStudy (ashish56):

g(6)=2*6-5=7 f[g(6)]=2*7+1=15